8. Consider the reaction: 2 H2S(g) 2 H2(g) +S2(g) Kc = 1.67 x 107 at 800°C What is the value of Kp at this temperature?
Answer #1

2H2sg = 2H2(g) + Sz(9) Ke = 1.67 X107 T= 800°c T=1073K on= 3-2 = ! Kp = kc (Rtjon Kp = 1.67x10 + (0.08221X1073) Kp = 1.47 X